IN one country applies the following for the national accounts: - Gross domestic product (GDP): 4900 - Income from abroad (net): 150 - Capital income from abroad (net): 170 - Transfers from abroad (net): -20 - Private consumption: 3580 - Investments including inventory investments: 950 - Public consumption: 1010 Calculate the current account!
Answer #1

Gdp= C+G+I+NX

4900= 3580 + 1010 + 950 + NX

4900 = 5540 + NX

NX= -640

Current account = NX + income from abroad + net transfer

= -640 + 150 - 20 = -510